Suppose that a main memory has 8192 blocks each block having 32 words. Show the memory address format (see figures 4.10 and 4.12 of your textbook for examples) for each of the cache organizations below.

a. a direct map cache with 64 lines of 32 words each

b. a set associative cache with 16 sets each with four lines of 32 words each

c. a fully associate cache with 128 lines of 32 words each

PLEASE EXPLAIN IN DETAIL
Answer #1

Number of blocks in Main memory size = 8192= 213

Block Size = 32 Words = 25 words

Memory Size = 213 * 25 = 218 words

Memory address = 18 bits

a) Direct Map Cache

Number of lines = 64 = 26 => Line offset = 6 bits

Line Size = 32 Words = 25 => Word offset = 5 bits

Remaining Tag bits = 18 - 5 - 6 = 7 bits

Tag Line Offset Word Offset
7 6 5

b) 4-Way Set-associative Cache

Number of Sets = 16 = 24 => Set offset = 4 bits

Line Size = 32 Words = 25 => Word offset = 5 bits

Remaining Tag bits = 18 - 5 - 4 = 9 bits

Tag SetOffset Word Offset
9 4 5

c) Fully-associative Cache

Number of Lines = 128 = 27

Line Size = 32 Words = 25 => Word offset = 5 bits

Remaining Tag bits = 18 - 5 = 13 bits

Fully associative cache only have tag and word offset as it directly maps main memory block.

Tag    Word Offset